OpenStudy (anonymous):

Find an equation of the circle with center C(7,-5) which is tangent to the x-axis

OpenStudy (anonymous):

......... well how big is the circle? ex. what's the radius

OpenStudy (anonymous):

it does not say what the radius is

OpenStudy (anonymous):

i know. ;)

OpenStudy (amistre64):

thats part of the solution process to figure it out. it gives you all the clues you need to determine the radius. What defines the radius of a circle?

OpenStudy (anonymous):

for the circle to touch the x axis, what does the radius have to be? try drawing this out. i like drawing things ^^

OpenStudy (amistre64):

might need to know what a tangent line to a circle represents as well

OpenStudy (anonymous):

is the radius 7 since it on the x-axis?

OpenStudy (amistre64):

good guess; but not quite, x axis is also the y=0 axis what is the distance from the center y value to the xaxis y value?

OpenStudy (anonymous):

im srry you just lost me

OpenStudy (amistre64):



OpenStudy (anonymous):

still confused

OpenStudy (amistre64):

what is the distance from center of the circle to the x axis? this measure will define the radius for us.

OpenStudy (amistre64):

the x axis IS y=0 the center of the circle has a y value of: y = -5 what is the distance between 0 and -5 ?

OpenStudy (anonymous):

-5?

OpenStudy (amistre64):

yes, but the negative part tells us a direction; its the 5 that tells us a distance. so lets define the radius as 5 units in length

OpenStudy (amistre64):

do you recall the general formula for the equation of a circle?

OpenStudy (anonymous):

yes it's (x-a)^2 + (y-b)^2=r^2

OpenStudy (amistre64):

good, and the center of the circle is defined as (a,b) and we just determined r; so can you show me how you think the solution would look?

OpenStudy (anonymous):

(x-7)^2+(y-(-5))^2=5

OpenStudy (anonymous):

5^2

OpenStudy (amistre64):

(a,b) = (7,-5); r=5 (x-a)^2 + (y-b)^2=r^2 (x-7)^2 + (y-(-5))^2 = 5^2 correct :)

OpenStudy (amistre64):

also acceptable is to write (y+5)^2 for the y part

OpenStudy (anonymous):

thxks! :)

OpenStudy (amistre64):

good luck ;)